OpenStudy (anonymous):

(x-3)(x+7)-(x+1)(x+5)=0

OpenStudy (anonymous):

\[ \mathrm{C.S.=-13} \]

OpenStudy (anonymous):

What part of this problem you don't understand exactly?

OpenStudy (anonymous):

The C.S. part

OpenStudy (anonymous):

\[(x-3)(x+7)-(x+1)(x+5)=0\\(x^2+4x-21)-(x^2+6x+5)=0\\-2x-26=0\\x=-13\]